A study on aggressive driving found that 52 % of drivers say that they have yelled at another driver. A random sample of 280 drivers in Halifax is selected and are asked whether they have yelled at other drivers.

Use the normal approximation to calculate the probability that there will be exactly 131 sampled drivers who will say yes. [Hint: use normal approximation]
Answer #1

Mean = np = 280*0.52 = 145.6

Standard deviation = = 8.36

Hence,

P(Exactly 131)

= P(130.5 < X < 131.5) [Continuity Correction]

= P(-1.81 < z < -1.69)

= 0.0104